# An Adaptive Computational Intelligence Approach to Personalised Health Assessment and Immune Age Characterisation from Common Haematological Markers

### v2

@@ Abstract
We introduce an approach to creating a simulated digital twin of the health of an individual based on a learning adaptive algorithm that learns the optimal reference values of the blood panel over time and associates an immune age score to compare with where if the biological age is lower than the score, an indication of wellness is provided. The score may also be useful for classification purposes. We demonstrate its efficacy against real and synthetic data from medically relevant cases, extreme cases, and empirical blood cell count data from 100K data records in the CDC NHANES survey that spans 13 years, from 2003 to 2016. We find that the score we introduce is informative when distinguishing healthy individuals from those with diseases, both self-reported and abnormal blood tests manifested, providing an entry-level score for patient triaging. We show that the score varies over time and is correlated with biological age, leading to the definition of an immune age as the inverse function of this relationship when different sets of analytes are taken only based on the results of the CBC test, providing clinical evidence of its potential relevance to the results of the CBC test, providing precision medicine and for personalised predictive healthcare.
Keywords: personalized health, predictive medicine, twin digital health, blood risk score, risk stratification, individual assessment, immune age
@@ 1 Introduction
Risk assessment tools are often needed to improve patient safety and patient outcomes. Risk assessment scores for patient sorting and triaging are often qualitative and are needed to select patient pathways and health monitoring choices across all levels of health care.
Risk scores are valuable predictive tools that, when implemented, can be utilised by clinicians with positive results, leading to better patient outcomes. Risk assessment measures for patient triage are often needed to optimise patient pathways and health monitoring across all levels of health care. However, most scores are defined for very specific conditions, such as heart disease, cardiovascular disease (CVD) or venous thromboembolism (VTE), to mention just a few. Others reflect factors such as risk of severe disease [23, 31], clinical acuity [10, 33], and long-term outcomes [30, 32]. For example, the Intermountain risk score [19] provides a complete / complete blood count (CBC / FBC) risk score shown to be associated with bleeding and in-hospital mortality based on red blood cell size, originally used as an anaemia predictor [26], and is often reported in FBC/CBC panels.
@@ 2 Methods
We will call this risk assessment score ”immune score”, given its connection to the immune system through blood cell subpopulation counts, in particular as registered in the most popular blood test, the Full Blood Count (FBC) or the Complete Blood Count (CBC).
@@ 2.1 Numerical score
This first formulation of the score takes the form of a linear function because it does not take into account possible interactions between analytes. We also assume that each analyte contributes equally to the immune score (this is likely to be only partially true, as some of the analytes are more or less medically informative than others, with this degree of informativeness itself varying with different conditions. Moreover, not all analytes are independent, and some analytes may be statistically dependent on others—suitable refinements will be introduced later).
For ease of comparison between successive versions of the score, all values are normalised within the range $[0,10]$. This means that each analyte will contribute with a maximum weight of
$w_{a}=\sqrt{\frac{100}{N}}.$
where $N$ is the number of analytes. From now on we will consider the example of $N=13$ typical for a CBC or FBC. From a geometrical point of view, a vector of values in the multidimensional immune space will be mapped to a point in a NND-“quadrant” going from 0 to $w_{a}$ that will be called the normalised immune space. This mapping will allow us to observe the diachronic evolution of an individual’s score as a trajectory in the normalised immune space where a notion of distance will come in handy.
@@
The adoption of a score within a maximum range of $[0,10]$ can be approached in different ways. One is to assume no theoretical maximum values for analytes and asymptotically approach the maximum score of 10, without ever reaching it in individual cases. We adopted an alternative approach, capping values to a pre-established “reasonable” maximum, beyond which all specific values mean the same. The pre-established maximum is two standard deviations beyond normal (healthy) ranges, which sets the statistical significance critical value alpha at 0.05 (i.e., 95% confidence interval).
@@
one-year old child, 2-6-year-old child, 6-12-year-old child}\}
@@
    Again, these values will come from a suitable set of cases. In the meantime, we will use the distance from the mean to either limit within the normal (healthy) range:
@@
  • The weighted vectors of each group (which will be used to normalise the possible values to a maximum immune score of 10):
@@
The norm of the normalised vector will be called the Normalised Immune Score, or NIS for short:
$|\bar{n}(\bar{c})$$|$.
Further adding to the novelty impact of the metric, the score also offers a way to incorporate medical knowledge in the form of weights (scalar or piecewise function) to be incorporated natively, though non-linear analytes can also be introduced as synthetic analytes that represent or simulate such linearities, and would do so in a clear mathematical fashion rather than in an obfuscatory statistical fashion, making it more transparent than dealing with non-linearity directly or using other possible statistical black-box approaches. Unlike statistical or machine learning approaches, the distance-metric-based score is transparent, so that clinicians can fully understand how it works, and even reproduces the calculation by hand with a justification at every step without loss of generality.
@@ 3 A numerical score derived from the CCIS
We expected the following property: If $v_{1},v_{2}$ are possible score values and $v_{1}<v_{2}$, it should be the case that $v_{2}$ indicates greater cause for concern than $v_{1}$.
On the other hand, values should be closely related to the colour codes. That is, values $v_{1}$, $v_{2}$ and $v_{3}$ correspond to CCIS scores of green, amber and red, respectively, and they should be ordered as follows:
$v_{1}<v_{2}<v_{3}.$
@@
In the following section, we present a procedure for translating the CCIS into a numerical value according to the above schema. We will call this value the NCCIS. The expected behaviour of the NCCIS with respect to the number of analytes and its deviation from the norm can be observed in Fig. 3.
Table 1: NHS reference values: Typical tabular presentation of a Complete/Full Blood Count test.
@@
Haemoglobin 152.3 115 165
Total White Cell Count 9.3 3.6 11.00
Platelet count 286.5 140 400
Neutrophils 6.9 1.8 7.5
Lymphocytes 3.6 1.0 4.0
Monocytes 0.3 0.2 0.8
Eosinophils 0.2 0.1 0.4
Basophils 0.2 0.02 0.1
Red cell count 5.5 3.8 5.8
Haematocrit 0.4 0.37 0.47
Mean Cell Volume (MCV) 93.3 80 100
Table 2: Mean values and standard deviations found in the NHANES 2003–2016 database for the 13 analytes of the immune scores.
@@
Hb 150.72 11.51 133.33 11.62
WBC 6.92 3.12 7.01 2.09
Plt 228 54.33 255.42 64.19
RBC 4.93 0.45 4.40 0.38
@@ 3.1 Calculating CCIS and NCCIS
Some of the values in the calculation of the NIS:
@@
Therefore the current NIS intervals for the different colours for individual analytes are:
@@
Figure 1: CCIS colour-based alerts. On the left, an individual with all values normal (healthy) and thus coded green. At the centre, an individual with two amber alerts. On the right-hand side, an individual with three red alerts, which produce a global red flag quantifying potential risk by colouring the numerical score depicted in the centre. In applications, colours following the guidelines of national health systems such as the NHS are recommended. For convenience, the numerical score can also be inverted to conform to the popular 0-10 grading scheme, where 10 is best and 0 is worst.
@@
The NCCIS is calculated using the following function:
$\mbox{NCCIS}(v)$ $=$ ${\sf if}\>\mbox{CCIS}(v)=\mbox{green }\>{\sf then}\>a_{green}(\mbox{NIS}(v))$
$\>{\sf else}\>{\sf if}\>\mbox{CCIS}(v)=\mbox{amber }\>{\sf then}\>a_{amber}(\mbox{NIS}(v))$
$\qquad\>{\sf else}\>a_{red}(\mbox{NIS}(v))$
where CCIS and NIS are functions calculating the respective scores and the functions $g_{a}$, $a_{a}$ and $r_{a}$ map the NIS to the intervals set in the table at the beginning of Section 4:
@@
and the NIS value is weighted according to a normalising weight:
@@
Table 3: Mapping of score segments and colours. For convenience, the numerical score can also be inverted to conform to the popular 0-10 grading scheme, where 10 is best and 0 is worst.
@@ 3.2 NHANES Age Prediction
The zero counts were removed from the NHANES data prior to all analysis. However, the sex was replaced by 1 for males and 2 to denote females, for the quantitative analysis. The sex groups were separated by health status as was determined by binning the immune scores and ages by the healthy range thresholds as indicated on Table 1, which are as follows: RBC (between 3.8 to 5.8), and for the multivariate measure, the healthy ranges were as follows: normal MCH (27-33), HCT (37-47), lymphocytes (1.0-4.0), and neutrophils (1.8-7.5). Above, or below these healthy count ranges, the individuals were binned into the unhealthy groups. Further, the error rates analysis for the age prediction between the actual ages and immune-score computed ages were performed using these binned subsets.
@@ 4 Results
Figure 2: Immune score space: The $x$ axis represents the deviation of all the analytes from the expected value, while the $y$ axis measures the total number of analytes. A negative standard deviation indicates that the value is lower than expected and a positive one indicates a larger than expected value. The cut-off seen on the negative $x$ axis shows the fact that real life values for analytes have a strict lower bound but, technically, have no upper bounds. For instance, it is not possible to have a negative white blood cell count.
Figure 4: Distribution of the Normalised Immune Score (NIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000–2018 database.The Healthy cohort is defined as those individuals with all analytes within the normal ranges as defined by the NHS (1). Others comprise the Unhealthy set. After filtering for adult status (18 years old or older) and completeness of analyte information, a total of 53,114 individuals are represented in the plot.
Figure 5: Distribution of the Normalised Colour-Coded Immune Score (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000–2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (1) and Unhealthy otherwise. After filtering for adult status (18 years old or older) and completeness of analyte information, a total of 53,114 individuals are represented in the plot.
@@
Fig. 4 shows how NIS values are distributed according to different (self-declared) conditions for individuals in the NHANES database. It must be noted that the self-reported conditions were not independently confirmed. Moreover, the survey participants did not distinguish between current or past diagnoses. This could be expected to introduce some noise, as some currently healthy people will be labelled as having a condition and some people with conditions will not have been diagnosed. We hope to improve accuracy by filtering the data or by adding other data sources in the future.
@@
Adult male with all values at the mean 0.00 Green 0.00
Adult male values slightly removed from mean within normal healthy reference interval 1.84 Green 1.84
Adult male with all abnormal values 10.00 Red 10.00
Adult male with leukocytosis and diabetes 4.15 Red 7.60
Adult male with pancytopenia 5.54 Red 8.18
Adult male with mycosis 4.69 Red 7.83
Adult female with fatigue 3.07 Amber 5.43
Adult female with shortness of breath 6.70 Red 8.68
Adult female with thrombocytopenia 5.50 Red 8.16
Adult (female) with infection 3.63 Red 7.39

### v4

# An Adaptive Computational Intelligence Approach to Personalised Health Assessment and Immune Age Characterisation from Common Hematological Markers

@@ Abstract
We introduce a simulated digital model that learns a person’s optimal blood health over time. Using a learning adaptive algorithm, our model provides a risk assessment score that compares an individual’s chronological age from birth to an estimation of a biological immune age derived from the score. We demonstrate its efficacy against real and synthetic data from medically relevant cases, extreme cases, and empirical blood cell count data from 100K data records in the Centers for Disease Control and Prevention’s National Health and Nutrition Examination Survey (CDC NHANES) that spans 13 years. We find that the score is informative when distinguishing healthy individuals from those with diseases, both self-reported and abnormal blood tests manifested, providing an entry-level score for patient triaging. The main contributions of this work is that the risk assessment score is not a machine learning black-box approach but a transparent adaptive learning algorithm and its calculation only requires existing common markers from a standard and most popular medical (and blood) test widely performed worldwide. This may allow a massively scalable approach to deploying a fast, personalised, sensitive and predictive derivative index without the need or application of a new test, assay, or prospective sampling as opposed to other biological ageing-related scores through a clinical platform and SaMD model.
Keywords: adaptive health learning, personalised health, predictive medicine, twin digital health, blood risk score, risk stratification, population assessment, individual assessment, immune age, biological age, senescence, longevity
@@ 1 Introduction
Risk assessment tools are often needed to improve patient safety and patient outcomes. Risk assessment scores used for initial patient classification or diagnosis and triaging (i.e., assessment of priority for healthcare services) are often qualitative and inter-subjective. Robust quantitative tools with clinical relevance are required to optimise health monitoring choices across all levels of healthcare systems.
Risk scores are valuable predictive tools that, when implemented, can be utilised by clinicians with positive results, leading to better patient outcomes. Risk assessment measures for patient triage are often needed to optimise patient pathways and health monitoring across all levels of health care. However, most scores are defined for very specific conditions, such as heart disease, cardiovascular disease (CVD) or venous thromboembolism (VTE), to mention just a few. Others reflect factors such as risk of severe disease [23, 31], clinical acuity [10, 33], and long-term outcomes [30, 32]. For example, the Intermountain risk score [19] provides a complete/full blood count (CBC/FBC) risk score shown to be associated with bleeding and in-hospital mortality based on red blood cell size, originally used as an anaemia predictor [26], and is often reported in FBC/CBC panels.
@@ 2 Methods
We will call this risk assessment score ”immune score”, given its connection to the immune system through blood cell subpopulation counts, in particular as registered in the most popular blood test, the FBC or the CBC.
@@ 2.1 Numerical score
This first formulation of the score takes the form of a linear function because it does not take into account possible interactions between analytes. CBC blood cell counts can interact through immune system responses. For example, an infection may increase neutrophils and decrease lymphocytes. RBC parameters can also be impacted by complex environmental cues such as nutrient levels, iron deficiency, stress conditions, or disorders. We also assume that each analyte contributes equally to the immune score (this is likely to be only partially true, as some of the analytes are more or less medically informative than others, with this degree of informativeness itself varying with different conditions. Moreover, not all analytes are independent, and some analytes may be statistically dependent on others—suitable refinements will be introduced later). Some examples of analytes that are independent of each other are RBC and WBC count (they do not directly influence the counts of each other). The counts of some analytes are dependent from each other. Examples include Neutrophils and WBC. In essence, neutrophils are a type of leukocyte (WBC), so the neutrophil count is included in and contributes to the total WBC. Another example is Hemoglobin and hematocrit. Hemoglobin measures the amount of hemoglobin protein in red blood cells. Hematocrit measures the percentage of blood volume taken up by red blood cells. Thus, Hematocrit is mathematically dependent on the hemoglobin concentration because one gives a subset or ratio of the other.
For ease of comparison between successive score versions, all values are normalised within the range $[0,10]$. Following user feedback, we then reverse the scale by subtracting the normalised value from 10. This means that each analyte will contribute with a maximum weight of
$w_{a}=\sqrt{\frac{100}{N}},$
where $N$ is the number of analytes. From now on we will consider the example of $N=13$ typical for a CBC or FBC. Before the scale inversion, from a geometrical point of view, a vector of values in the multidimensional immune space will be mapped to a point in a NND-“quadrant” going from 0 to $w_{a}$ that will be called the normalised immune space. This mapping will allow us to observe the evolution of an individual’s score as a trajectory in the normalised immune space where a notion of distance will come in handy.
@@
The adoption of a score within a maximum range of $[0,10]$ can be approached in different ways. One is to assume no theoretical maximum values for analytes and asymptotically approach the minimum score of 0, without ever reaching it in individual cases. We adopted an alternative approach, capping values to a pre-established “reasonable” maximum, beyond which all specific values mean the same. The pre-established maximum is two standard deviations beyond normal (healthy) ranges, which sets the statistical significance critical value alpha at 0.05 (i.e., 95% confidence interval).
@@
one-year old child, 2-6-year-old child, 6-12-year-old child, 12-18-year-old teenager}\}
@@
    Again, these values will be empirically calculated. In the meantime, we will use the distance from the mean to either limit within the normal (healthy) range:
@@
  • The weighted vectors of each group (which will be used to normalise the possible values to a maximum immune score, before reversion, of 10):
@@
The value obtained by subtracting the normalised vector $\bar{n}(\bar{c})$ from 10 is defined as the Normalised Immune Score, abbreviated as NIS:
$\text{NIS}=10-\|\bar{n}(\bar{c})\|.$
We will denote the uninverted score by $\text{NIS}^{*}$, which is defined as $\text{NIS}^{*}=\|\bar{n}(\bar{c})\|$.
One novel aspect of this metric is the ability to incorporate medical knowledge into the score calculation using weights (scalar or piece-wise functions). This allows adding clinical expertise to the data-driven score in a transparent mathematical way, rather than using opaque statistical black-box approaches. Statistical black-box approaches refer to complex statistical modelling techniques that are not easily interpretable (e.g., neural networks). The weights make it possible to adjust analyte contributions in a nonlinear fashion if needed. Synthetic analytes can also be introduced to simulate nonlinear relationships in a clear mathematical form, rather than dealing with nonlinearities directly in a statistically obscure way. Unlike statistical or machine learning approaches, the distance-metric-based score is transparent, so that clinicians can fully understand how it works, and even reproduces the calculation by hand with a justification at every step without loss of generality.
@@ 2.2 Colour-based score
Keeping the range to within two standard deviations when assessing a score or metric has the following statistical implications: 1) About 95% of observations fall within 2 standard deviations of the mean in a normal distribution. Thus, restricting the range cuts off the most extreme 5% of outliers. This makes the statistics more robust by removing very unusual values (outliers) that could skew results. But it also reduces the variability seen in the data. However, it may also provide an overly narrow or skewed view of the true variability in the evaluated metrics. For the goals of our analysis, i.e., sample CBC analysis, assuming a normal distribution justifies this rationale and its implications.
@@ 3 A numerical score derived from the CCIS
We expected the following property: If $v_{1},v_{2}$ are possible score values and $v_{1}>v_{2}$, it should be the case that $v_{2}$ indicates greater cause for concern than $v_{1}$.
On the other hand, values should be closely related to the colour codes. That is, values $v_{1}$, $v_{2}$ and $v_{3}$ correspond to CCIS scores of green, amber, and red, respectively, and they should be ordered as follows:
$v_{1}>v_{2}>v_{3}.$
@@
In the following section, we present a procedure for translating the CCIS into a numerical value according to the above schema. We will call this value the NCCIS. The expected behaviour of the NCCIS concerning the number of analytes and its deviation from the norm can be observed in Fig. 3.
Table 1: NHS reference values: Typical tabular presentation for illustration purposes of a Complete/Full Blood Count test. Minimum and maximum values show minimum and maximum between male/female values. Typical values used in the U.S. are available at https://myhematology.com/red-blood-cells/full-blood-count-and-other-hematology-reference-ranges/
@@
Haemoglobin (HgB) 152.3 g/L 115 g/L 180 g/L
Total White Cell Count (WBC) 9.3 g/L 3.6 g/L 11.00 g/L
[…]
Mean Cell Volume (MCV) 93.3 fL 80 fL 100 fL
Haematocrit 0.4 L/L 0.37 L/L 0.47 L/L
MC Haemoglobin (MCH) 29 pg 27 pg 32 pg
MCH Concentration (MCHC) 315 g/L 310 g/L 350 g/L
[…]
Table 2: Mean values and standard deviations found in the NHANES 2003–2016 database for the 13 analytes of the immune scores (same units as above).
@@
Haemoglobin (HgB) 150.72 11.51 133.33 11.62
Total White Cells 6.92 3.12 7.01 2.09
Platelet count (PLT) 228 54.33 255.42 64.19
Total Red Cells (RBC) 4.93 0.45 4.40 0.38
@@ 3.1 Calculating CCIS and NCCIS
The computation of NCCIS is performed using the uninverted $\text{NIS}^{*}$. Some of the values used in the calculation of the $\text{NIS}^{*}$ include:
@@
Therefore the current $\text{NIS}^{*}$ intervals for the different colours for individual analytes are:
@@
Figure 1: CCIS colour-based alerts. On the left, an individual with all values normal (healthy) and thus coded green. At the centre, an individual with two amber alerts. On the right-hand side, an individual with three red alerts, which produce a global red flag quantifying potential risk by colouring the numerical score depicted in the centre. In applications, colours following the guidelines of national health systems such as the NHS are recommended. For convenience, the numerical score has been inverted (with respect to $\text{NIS}^{*}$) to conform to the popular 0-10 grading scheme, where 10 is best and 0 is worst.
@@
The $\text{NCCIS}^{*}$ is calculated using the following function:
$\mbox{NCCIS}^{*}(v)$ $=$ ${\sf if}\>\mbox{CCIS}(v)=\mbox{green }\>{\sf then}\>a_{green}(\mbox{NIS}^{*}(v))$
$\>{\sf else}\>{\sf if}\>\mbox{CCIS}(v)=\mbox{amber }\>{\sf then}\>a_{amber}(\mbox{NIS}^{*}(v))$
$\qquad\>{\sf else}\>a_{red}(\mbox{NIS}^{*}(v))$
where CCIS and $\text{NIS}^{*}$ are functions calculating the respective scores and the functions $g_{a}$, $a_{a}$ and $r_{a}$ map the $\text{NIS}^{*}$ to the intervals set in the table at the beginning of Section 4:
@@
and the $\text{NIS}^{*}$ value is weighted according to a normalising weight:
@@
Finally, as in the case of NIS, we invert the scale to obtain the NCCIS:
$\text{NCCIS}=10-\text{NCCIS}^{*}.$
Table 3: Mapping of score segments and colours. For convenience, the numerical score has been inverted to conform to the popular 0-10 grading scheme, where 10 is best and 0 is worst.
@@ 3.2 NHANES Age Prediction
The zero counts were removed from the NHANES data prior to all analysis. However, the sex was replaced by 1 for males and 2 to denote females, for the quantitative analysis. The sex groups were separated by health status as was determined by binning the immune scores and ages by the healthy range thresholds as indicated on Table 1, which are as follows: Red Blood Cell (RBC) count (between 3.8 to 5.8), and for the multivariate measure, the healthy ranges were as follows: normal Mean Corpuscular Hemoglobin (MCH) (27-33), Hematocrit (HCT) (37-47), lymphocytes (LYMP) (1.0-4.0), and neutrophils (NEUT) (1.8-7.5). Above, or below these healthy count ranges, the individuals were binned into the unhealthy groups. Further, the error rates analysis for the age prediction between the actual ages and immune-score computed ages were performed using these binned subsets.
Figure 2: Immune score space: The $x$ axis represents the deviation of all the analytes from the expected value, while the $y$ axis measures the total number of analytes. A negative standard deviation indicates that the value is lower than expected and a positive one indicates a larger than expected value. The cut-off seen on the negative $x$ axis shows the fact that real-life values for analytes have a strict lower bound but, technically, have no upper bounds. For instance, it is not possible to have a negative white blood cell count.
Comparison Sample Size (n) p-value
Healthy vs Unhealthy 7007 vs 6074 P ¡ 0.0001
Self rep. Healthy vs Unhealthy 11976 vs 6074 P ¡ 0.0001
Healthy vs Anemia 7007 vs 980 P ¡ 0.0001
Healthy vs HIV 7007 vs 62 P ¡ 0.0001
Healthy vs Hodgkin’s Disease 7007 vs 37 P ¡ 0.0001
Healthy vs Leukemia 7007 vs 18 P ¡ 0.0001
Healthy vs Blood Cancer 7007 vs 8 P ¡ 0.0001
Table 4: Differences in NIS Mean Values. All comparisons were performed using ANOVA.
Figure 4: Distribution of the Normalised Immune Score (NIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000–2014 database. The Healthy cohort is defined as those individuals with all analytes within the normal ranges as defined by the NHS (1), others comprise the Unhealthy set. The Self rep. healthy cohort compromises individuals that self-reported health of 3 (”Good”) or less The number of individuals in each cohort are shown in table 4.
Comparison Sample Size (n) p-value
Healthy vs Unhealthy 7007 vs 6074 P ¡ 0.0001
Self rep. Healthy vs Unhealthy 11976 vs 6074 P ¡ 0.0001
Healthy vs Anemia 7007 vs 980 P ¡ 0.0001
Healthy vs HIV 7007 vs 62 P ¡ 0.0001
Healthy vs Hodgkin’s Disease 7007 vs 37 P ¡ 0.0001
Healthy vs Leukemia 7007 vs 18 P ¡ 0.0001
Healthy vs Blood Cancer 7007 vs 8 P ¡ 0.0001
Table 5: Differences in NCCIS Mean Values. All comparisons were performed using ANOVA.
Figure 5: Distribution of the Normalised Colour-Coded Immune Score (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000–2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (1) and Unhealthy otherwise. Self rep. healthy compromises individuals that self-reported health of 3 (“Good”) or less. The number of individuals in eah cohort is shown in table 5. One explanation for the bimodality of some of these plots is the nature of the source data as self-reported where respondents may or may not have been ill at the time of giving their answers, or from the application of the separating cutoff value from NIS. In either case, both score-separating capabilities are highly statistically significant.
@@ 4 Results
Figure 4 illustrates the distribution of NIS values across various (self-declared) conditions for individuals in the NHANES database. The table 4 presents the number of data points used, along with the respective ANOVA results to test statistical significance. It must be noted that the self-reported conditions were not independently confirmed. Moreover, the survey participants did not distinguish between current or past diagnoses. This could be expected to introduce some noise, as some currently healthy people will be labelled as having a condition and some people with conditions will not have been diagnosed. We hope to improve accuracy by filtering the data or by adding other data sources in the future.
@@
Adult male with all values at the mean 10.00 Green 10.00
Adult male values slightly removed from mean within normal healthy reference interval 8.16 Green 8.16
Adult male with all abnormal values 0.00 Red 0.00
Adult male with leukocytosis and diabetes 5.85 Red 2.40
Adult male with pancytopenia 4.46 Red 1.82
Adult male with mycosis 5.31 Red 2.17
Adult female with fatigue 6.93 Amber 4.57
Adult female with shortness of breath 3.30 Red 1.32
Adult female with thrombocytopenia 4.50 Red 1.84
Adult (female) with infection 6.37 Red 2.61